zepdrix (zepdrix):

Fun math problem. This was a cool little problem I saw at work today. When the time is 8:30, what acute angle do the hands on a clock make? (In degrees)

OpenStudy (charlotte123):

33? o-o

OpenStudy (shamil98):

30 degrees?

zepdrix (zepdrix):

Hmmmm, nope :x

OpenStudy (charlotte123):

IHTS A TRICK QUESTION

zepdrix (zepdrix):

There is a small tricky part to it:) But's not a trick question lol.

OpenStudy (mony01):

maybe 10 degrees

OpenStudy (charlotte123):

45 degrees? :)

OpenStudy (shamil98):



OpenStudy (anonymous):

60 degrees?

OpenStudy (anonymous):

75

zepdrix (zepdrix):

Euler for the winnnn! \c:/

OpenStudy (anonymous):

yay :)

OpenStudy (charlotte123):

HUZZAH

OpenStudy (shamil98):

o.o

OpenStudy (charlotte123):

┌(-_-) ┐└(-_-)┘┌(-_-) ┐└(-_-)┘ ♪┏(・o・)┛♪┗ ( ・o・) ┓♪┏ ( ) ┛♪┗ (・o・ ) ┓♪┏(・o・)┛

OpenStudy (mony01):

75 degrees

OpenStudy (charlotte123):

XD

OpenStudy (anonymous):

How is it 75?

zepdrix (zepdrix):

60 is along the right line of thinking, that was my guess when I first saw the problem. You have to remember, when it's 8:30, the hour hand is NOT on the 8. It's halfway between the 8 and 9

OpenStudy (charlotte123):

http://www.geekinterview.com/question_details/13081 o-o

OpenStudy (charlotte123):

well den .-. XD

OpenStudy (anonymous):

my math was: \[\left( \frac{ 3 }{ 12 } -\frac{ 1 }{ 2 }*\frac{ 1 }{ 12 }\right)360°\]

OpenStudy (shamil98):

forgot its a circle so its out of 360 degrees ._. and 1/4 of it is 90

OpenStudy (anonymous):

hahaha, that was funny.

OpenStudy (shamil98):

was thinking of a triangle lul

zepdrix (zepdrix):

ya fun stuff c: hah

zepdrix (zepdrix):

triangle? XD interesting

OpenStudy (charlotte123):

out of all things, A TRIANGLE XD @zepdrix

OpenStudy (charlotte123):

shamil XD

OpenStudy (shamil98):

Brain farts

zepdrix (zepdrix):

Are most clocks shaped like a triangle? :3 lolol

OpenStudy (charlotte123):

eat ur peaches

OpenStudy (shamil98):

in a parallel universe yes

OpenStudy (charlotte123):

@zepdrix looks like shamils is o-o dont blame the poor guy :( LAWWWWWWWWWWWWLLLLLLLLLLLLLL

zepdrix (zepdrix):

lol :3

OpenStudy (charlotte123):

this was fun :D thank u zep ლ(́◉◞౪◟◉‵ლ)